OpenStudy (anonymous):

find the value of y.............. cos y = 8/13

OpenStudy (anonymous):

52.02

OpenStudy (anonymous):

Cos y =8/13 y= Cos^-1(8/13)

OpenStudy (anonymous):

thanx you guys

OpenStudy (phi):

type acos(8/13) in degrees = into the google search window.